A 3-phase, 380 V, 50 Hz , 6 pole, Y-connected induction motor has R1 = 0.1 W , X1 = 0.4 W, R2’ = 0.12 W , X2’ = 0.4 W, Xm = 20 W , Pr = 1 kW. The motor is operated at 970 rpm what is the The ratio of the stator starting current to full-load current (Istart / Ifull-load)?
Answer #1

[page-1. lalui Given, 3-phase, Y-connected induction niotor: 380V, 50 Hz, 6 pole R, = 0.14 x = 0.42 R2:20.12 , x2 = 0.41 Xm=2Ist oil 30-4 > (page 2 1 2q = (0.114 tjo.391) + 219.39 IV KUL) By KULI Ist = 219.393 (0.1tjo.4+0.114 14 tj0:39 Ist = 219.393Zeqi 그 j 20+4+jot > Zeqi [page-37 (jro) (4+jo.4) (20290) (4.019/5.5 (4+j20-4) (20290*) (4.01975-11) (20.788 178.9) = 3.8666 [